OpenStudy (steve816):

Someone explain this to me! How is cos(90) in radians a negative number?

jimthompson5910 (jim_thompson5910):

cos(90 degrees) = 1 cos(90 radians) = -0.44807361612918 keep in mind that `180 degrees = pi radians` so we can use this to convert 90 radians = (90 radians) * (180 degrees/pi radians) = 90*(180/pi) = 5,156.6201561774 degrees

jimthompson5910 (jim_thompson5910):

90 radians = 5,156.6201561774 degrees then find the coterminal angle to `5,156.6201561774 degrees` that is between 0 and 360